After three months, $8200 deposited in a savings account with simple interest had grown to a total of $8364. What was the interest rate?
Answer by Boreal(15235)

I=prt
p=8200
r unknown
t=1/4
i=164
164=8200*1/4*r=2050r
r=164/2050=0.08 or 8%